# LLMs Reproduce Stereotypes of Sexual and Gender Minorities

### v1

@@ 4.4 LLM Responses to SCM Prompts
Table 4: Most frequent keywords for each group obtained by LLM elicitation. Words in orange associate with Warmth and words in blue with Competence. Words in italics are used by multiple models for the same group while words in bold were frequently elicited from human survey participants for the same group (Table 3).
Across both human and LLM survey results, there is a persistent pattern of associating nonbinary people, and often bisexuals, with low rates of Warmth and Competence. By contrast, women are strongly associated with Warmth, men with Competence, and heterosexuals with either or both. GPT strongly follows the patterns of the survey responses, as does LLaMA to a lesser degree. Gemini is a clear outlier, often skewing strongly in favor of Warmth or Competence, but not in a consistent way.
@@ 5 SCM Axes in LLM Text Generation
While surveys enable us to assess whether LLMs reflect societal stereotypes in a way consistent with social psychology findings, they don’t model likely use cases, and thus don’t fully establish that a user might encounter such behavior. Story generation is a common method of testing bias in NLG models (e.g., Lucy and Bamman (2021), Narayanan Venkit et al. (2023), Bai et al. (2024), and Kumar et al. (2024)), with a focus often placed on how characters are described. Indeed as we note in the introduction, LLM providers widely promote LLMs as writing aids. Since we have established that SCM axes can be used to measure LLM bias in the controlled setting of a survey, we now turn to their use in measuring bias in this more realistic setting.
@@ 5.1 Results
Since text generation is much more open-ended than the elicitation task, we need a way to focus on descriptive words. We used SpaCy Honnibal and Montani (2017) to identify nouns, adjectives, and verbs from the generated texts, focusing solely on these words in our subsequent analysis. Rates of Warmth and Competence-coded terms—coded using the methods of Section 4.2—appear in Figure 8. Since generation produces far greater and more varied words than elicitation, percentages are lower relative to the survey. They are also not as consistent with the human survey as earlier results, but we nevertheless observe that the relative associations of Competence between groups generally track those of the survey. They are also consistent with the survey in the sense that they prefer Warmth to Competence for most groups.
@@
Figure 8: Percentage of Warmth and Competence terms obtained by LLM text generation (cf. Figure 4). Since terms unrelated to these concepts occur in the data, percentages do not sum to 100.
@@
To understand the results qualitatively, we first looked at the most frequent words for each group. Consistently across all LLMs for all groups, these tended to be generic, associated with bodily descriptions (e.g., “eyes”, “hair”) and location (e.g., “village”, “town”). “Love” was also a common term, particularly for sexual orientation groups; indeed it is the most frequently generated word for all those groups in GPT and Gemini. The clearest difference was the high frequency of the group-specific term (e.g., “man”, “woman”, “lesbian”) in the output for that group, as we might expect. This suggests that the textual output is often somewhat formulaic, with similar structural elements.
To better understand the qualitative differences in results between texts about each group, it is helpful to focus on words that the models specifically associate this group. Inspired by Wan et al. (2023), we use the Odds Ratio for this. This quantity is the ratio of two conditional probabilities: that of generating a word, conditioned on the group of interest, and that of generating the same word conditioned on all other groups. Table 6 shows the highest associated words for each group based on their Odds Ratio for that group, highlighting words coded for Warmth and Competence.
@@
Qualitatively, many of the generated passages reinforce the Warmth-Competence stereotypes found throughout our results. For example, stories about women focus on kindness, compassion, and love (e.g. “ability to heal others", “a passionate advocate for social justice"). Those about men focus on knowledge, courage, and adventure (e.g. a man learning to rock climb grows “stronger and more confident"). A nonbinary person “often felt misunderstood" by others and “whispers and sideways glances" followed them. A lesbian faces “discrimination and marginalization" throughout their life. A bisexual is “condemned" and called “a deviant, a threat to the village’s morals." These passages emphasize marginalization and pain for minoritized identities, which, while likely representative of their data, also present a limited representation of individuals from these groups and reifies painful experience as most representative of their lives even in creative stories. Indeed, all LLMs frequently generated words suggesting struggle (e.g., “challenges”, “justice”, “messy”) in texts about nonbinary people, bisexuals, and lesbians, a pattern recently found by Dhingra et al. (2023) as well.
6 Discussion
This paper attempts to synthesize two distinct threads in the research on gender bias in NLP. The first thread, exemplified by Dhingra et al. (2023), aims to move the discussion of gender past a binary distinction of men and women, a gap pointed out by (Stanczak and Augenstein, 2021) and (Devinney et al., 2022) that dovetails with other efforts to include queer experiences in the scope of NLP research (Lissak et al., 2024, e.g.). The second thread aims to move measurement of bias towards a surer footing by articulating harms (Blodgett et al., 2020) and operationalizations (Goldfarb-Tarrant et al., 2023). To do this, we ground our measurement of stereotypes in the Stereotype Content Model (Fiske et al., 2002), a well-studied theory of social psychology which has been empirically shown to correlate with emotions and behaviours towards different groups (Cuddy et al., 2007).
It is important to acknowledge that both of these threads continue to evolve, and the methods we use here are only a step in the right direction. For example, we use sets of keywords (Table 4) from SCM studies which were validated at the time of those studies as reasonable operationalizations of their respective concepts. But these concepts, their methods of validation, and the results of those validations can change over time. Indeed, there are recent efforts to review and update these operationalizations (Halkias and Diamantopoulos, 2020; Friehs et al., 2022) which partly affirm and partly question their value. Still, when we have re-analyzed our data from the rating task (Figure 2 and Section 4.4) using only those words with stronger validity according to Halkias and Diamantopoulos (2020), we found little change in the relationship between human and LLM ratings and, more generally, in overall patterns of results. In another direction, there is an expanded set of concept operationalizations devised by Nicolas et al. (2021), which we use in our experiments to automatically classify words as coded for Warmth and Competence. These conceptualizations have been used for human studies (Nicolas et al., 2022), and, concurrently with our work, for studies of LLMs (Nicolas and Caliskan, 2024). The latter, though it includes the groups that we focus on here, focuses on a broad analysis of many different groups, rather than the targeted and deep analysis that we present here.
@@ 7 Conclusion
This paper presents a test of three large language models—GPT 3.5, Gemini 1.5, and LLaMA 2—on gender and sexual orientation stereotypes by utilizing the Stereotype Content Model. The models were tested in two stages, first with explicit request for stereotypical perceptions and second in the application of creative text generation. A survey was run with which to compare LLM results, a necessary addition given the absence of some groups in past work.
@@
Certain themes of Warmth and Competence perceptions for the groups of study were clearly evident in the LLMs. Consistent with prior work on human perceptions of stereotype, women were strongly associated with Warmth and men with Competence. These results help frame our novel study of both human and LLM stereotypes of sexual and gender minorities. All of the minoritized groups that we study—gay men, lesbians, bisexuals, and nonbinary people—were rated consistently lower on Competence, with the most powerless of these—bisexuals and nonbinary people—also rating consistently lower on Warmth than most other groups. Heterosexuals, in contrast, were associated with normalcy, and often rated more highly by both people and LLMs on both axes. The same patterns were still evident in text generation, though more subtly in measurement. This result—along with the alignment with survey responses—further legitimates the SCM as a helpful measure of stereotype, though further work remains to be done.
@@
We found some differences in the behavior of the LLMs: GPT most accords with survey participants throughout testing, with LLaMA close as well. Gemini diverges the most from the survey responses but many of the themes still hold.
@@
LLM vendors continue to promote their products as creative writing assistants. Consistent with many other studies on bias in NLP, we emphasize that these tools can amplify biases towards sexual and gender minorities, a diverse group which has received relatively little attention in the research literature. We urge users of these tools to gain awareness of these types of risks and to exercise caution when using these tools as advertised.
@@ Acknowledgments
We thank Fengyu Liu and Yuanqi Shi for helpful discussion of this work; and Sharon Goldwater, Coleman Haley, Oli Liu, Yen Meng, and Sung-Lin Yeh for comments on earlier drafts of this paper.

### v2

@@ 4.4 Results
Table 3: Most frequent keywords obtained by Elicitation from humans and LLMs. Words in orange associate with Warmth and words in blue with Competence using the method in Section 4.2. Words in italics are used by multiple models for the same group. Words in bold were the most frequent words elicited from both humans and at least one LLM for the same group. Groups are ordered by ascending human rating for Competence (Figure 2).
Elicitation
Table 3 and Figure 4 summarize the results. We again see similar patterns to the other survey questions: terms elicited for women associate with Warmth (e.g. “nurturing”); for men with Competence (e.g. “leader”); for heterosexuals with both, and, more qualitatively, with normalcy (e.g. “normal” and “natural”). In contrast, nonbinary and bisexual people elicit more negative terms, including words relating to confusion (e.g. “confused”, “lost”, “indecisive”). Keywords for the minority groups include “courageous”, “brave”, “strength” and “resilient”, which are coded for Competence but also allude to historical discrimination.
10101010202020203030303040404040Humanof termsCompetenceWarmth
00252525255050505075757575100100100100GPTof terms
00252525255050505075757575100100100100Geminiof terms
Figure 4: Percentages of Warmth and Competence terms obtained by Elicitation, using the method of Section 4.2 to associate words to concepts. Since unrelated terms can occur, percentages do not sum to 100. Groups ordered by human rating for Competence (Figure 2).
GPT and LLaMA tend to follow the patterns of human ratings, with GPT following them quite closely. Gemini is again skewed, producing either Warmth or Competence for each group; but its skew differs from the one observed in selection, where the preferred category differs across tasks for nonbinary people, bisexuals, and heterosexuals.
Critically, the different survey questions yield fairly consistent results: they recapitulate SCM findings about stereotypes of men, women, and gay men (Fiske et al., 2002), and repeatedly elicit a perception of more negative stereotypes of sexual and gender minorities, most strongly of nonbinary and bisexual people.
@@ 5 SCM Axes in LLM Text Generation
Surveys assess whether LLMs reflect societal stereotypes in a way consistent with social psychology findings. But they don’t model likely use cases, so they don’t show that LLM users might encounter stereotypes. As we note in the introduction, LLM providers promote LLMs as creative writing tools. Story generation is often use to test for bias (e.g. Lucy and Bamman, 2021; Narayanan Venkit et al., 2023; Bai et al., 2024; Kumar et al., 2024), focusing on how characters are described, so we now adapt the SCM concepts measure stereotypes in this more realistic setting.
@@ 5.1 Results
Text generation is more open-ended than even the elicitation task, so we need a way to focus on descriptive words. We used SpaCy Honnibal and Montani (2017) to identify nouns, adjectives, and verbs in generated texts, focusing solely on these words in subsequent analyses. Quantitative results (Figure 5) are not as consistent with the human survey as earlier results, but we still observe that the relative associations of Competence track those of the survey. Results are also consistent with the survey in the sense that they prefer Warmth to Competence for most groups.
@@
Figure 5: Percentage of Warmth and Competence terms obtained by LLM text generation, using the method of Section 4.2 to associate words to concepts. Since unrelated terms occur, percentages do not sum to 100.
@@
To understand the results qualitatively, we first looked at the most frequent words for each group. Consistently across all LLMs and groups, these tended to be generic: bodily descriptions (e.g., “eyes”, “hair”) and locations (e.g., “village”, “town”). “Love” was also a common term, particularly for gay men, lesbians, and bisexuals; indeed it is the most frequent word that GPT and Gemini generate for all of these groups . Group-specific terms (e.g., “man”, “woman”, “lesbian”) also appear frequently in the output for that group. These observations suggest that generated texts are formulaic, with similar structures.
@@
To focus on words that the models strongly associate with each group, we borrow an idea from Wan et al. (2023), and compute the Odds Ratio (OR) for each word and group. OR is the ratio of two conditional probabilities: that of generating a word conditioned on the group of interest, and that of generating the same word, not conditioned on any group. The results (Table 5) show that generated texts reinforce the Warmth-Competence stereotypes found throughout our results. For example, stories about women focus on Warmth (e.g. “ability to heal others", “a passionate advocate for social justice"); and about men on Competence (e.g. a man learning to rock climb grows “stronger and more confident"). A nonbinary person “often felt misunderstood" by others and “whispers and sideways glances" followed them. A lesbian faces “discrimination and marginalization" throughout their life. A bisexual is “condemned" and called “a deviant, a threat to the village’s morals." These passages emphasize marginalization and pain for minoritized identities, reifying painful experience as most representative of their lives even in creative stories. Indeed, all LLMs frequently generated words about struggle (e.g., “challenges”, “justice”, “messy”) for nonbinary people, bisexuals, and lesbians, a pattern also observed by Dhingra et al. (2023).
@@ 6 Conclusion
We’ve attempted to synthesize two distinct threads in the research on gender bias in NLP. The first, exemplified by Dhingra et al. (2023), aims to move the discussion of gender past a binary distinction of men and women, dovetailing with other efforts to include queer experiences in the scope of NLP research (Lissak et al., 2024, e.g.). The second aims to move measurement of bias towards a surer footing by articulating harms (Blodgett et al., 2020) and operationalizations (Goldfarb-Tarrant et al., 2023). To do this, we ground our measurement of stereotypes in the SCM (Fiske et al., 2002), a well-studied theory of social psychology which has been empirically shown to correlate with emotions and behaviours towards different groups (Cuddy et al., 2007).
Using the SCM, we tested three large language models—GPT 3.5, Gemini 1.5, and LLaMA 2—for stereotypes towards gender and sexual orientation minorities. Following Jeoung et al. (2023) we tested the models just as one might test a human subject in a psychology experiment. To compare with real societal stereotypes, we ran a parallel study with human subjects. Where our experiments overlap with previous research, they are consistent. But our human survey contains focuses on sexual and gender minorities not included in past surveys, and we analyze the data in more detail than past surveys of either humans or LLMs, which have tended to be broad. These novel results help us to understand the specific ways in which stereotypes of these groups are perpetuated by LLMs.
@@
All of the minoritized groups that we study—gay men, lesbians, bisexuals, and nonbinary people—were rated consistently lower on Competence, with the most powerless of these—bisexuals and nonbinary people—also rating consistently lower on Warmth than most other groups. Heterosexuals, in contrast, were associated with normalcy, and often rated more highly by both people and LLMs on both axes. These patterns persist in text generation: though the quantitative results are more subtle, qualitative results demonstrate starkly stereotyped portrayals of different groups.
@@
We observed some differences in the behavior of the LLMs: GPT mostly accords with survey participants throughout testing, with LLaMA also similar. Gemini diverges the most from the survey responses but many themes still hold.
@@
LLM vendors continue to promote their products as creative writing assistants. Consistent with other studies on bias in NLP, we emphasize that these tools amplify biases towards sexual and gender minorities, a diverse group which has received relatively little attention in the research literature. We urge LLM users to gain awareness of these risks and to exercise caution when using them as advertised.
Limitations
Our study considers attitudes exhibited towards specific social groups in the English language, and makes no claim about how the results might change when considering other languages.
We screened human participants for English fluency, but not location or country of origin. Attitudes towards different social groups may vary across countries and geographic regions, and since our analysis is aggregate, it may not fully reflect this complexity. Similarly, the LLMs we study were trained on very large datasets whose details are not publicly known, but which likely contain examples from many different English-speaking countries and geographic regions. Hence, we cannot know how closely the mix of Englishes represented in training data of the LLMs reflects the mix of Englishes represented in our human survey.
Our survey does not ask participants about their personal opinion. This is the established approach to SCM survey methods Fiske et al. (2002); Cuddy et al. (2007), and makes our results more amenable to comparisons with that work. However, it is known that people’s perceptions may be biased by exposure to news coverage. For example, after a rare disaster occurs (such as a plane crash) survey respondents are likely to over-estimate the risk of such a disaster. Similarly, it’s possible that in the political climate of summer 2024, when we conducted our survey, respondents have a biased perspective of how society views sexual and gender minorities, based on widespread and polarized political news coverage in several English-speaking countries. In contrast, the training data of the LLMs we study likely almagamates attitudes across many years, rather than at a specific point in time.
Ethical Considerations
Our study involved human participants. We obtained approval from the University of Edinburgh School of Informatics Ethics Committee (application 783573) on June 5th, 2024, prior to commencing any work with participants. Since participants were asked to reflect on stereotypes about groups that they may be members of, we were aware that this may cause distress. Participants were advised of this in advance, and directed to mental health resources in the case of distress. Participants were also advised that they could withdraw from the study at any time without explanation.
@@ Acknowledgments
We thank Fengyu Liu and Yuanqi Shi for helpful discussion of this work; and Sharon Goldwater, Coleman Haley, Oli Liu, Yen Meng, Sung-Lin Yeh, and anonymous TACL reviewers for comments on earlier drafts of this paper.